OpenStudy (anonymous):

Solving exponential functions: 8x=12,143

OpenStudy (anonymous):

I'm not sure where to start.

OpenStudy (anonymous):

I got 0

OpenStudy (anonymous):

can you show me how you did it please?

OpenStudy (anonymous):

No its not 0

OpenStudy (campbell_st):

so the question is \[8^x = 12143\]

OpenStudy (campbell_st):

if it is... take the log of both sides \[\ln(8^x) = \ln(12143)\] using log laws it can be written as \[xln(8) = \ln(12143)\] so \[x = \frac{\ln(12143)}{\ln(8)}\] just evaluate for the solution.